### v1

# Current Cross-Correlation Spectroscopy of Majorana Bound States

Abstract
The clock speed of topological quantum computers based on Majorana zero mode (MZM)-supporting nanoscale devices is determined by the time taken for electrons to traverse the device. We employ the time-dependent Landauer-Büttiker transport theory for current cross-lead correlations in a superconducting nanowire junction hosting MZMs. From the time-dependent quantum noise, we are able to extract traversal times for electrons crossing the system. After demonstrating a linear scaling of traversal times with nanowire length, we present a heuristic formula for the traversal times which accurately captures their behaviour. We then connect our framework to a proposed experimental verification of this discriminant between spurious and genuine MZMs utilizing time-resolved transport measurements.
Introduction.— Modern topological computing protocols are increasingly dominated by the exchange and manipulation of non-local, zero-energy, zero-charge, zero-spin quasiparticles known as Majorana Zero Modes (MZMs) [1, 2, 3]. Recent experimental work has demonstrated the ability to engineer MZMs in superconducting heterostructures, particularly those based on InAs or InSb [4, 5, 6, 7], and more recently Ge [8].
Among the most promising systems for realizing MZMs in the laboratory are proximitized nanowires [9, 4, 10, 11, 12, 13, 14, 15, 16], an example of which is shown schematically in Fig. 1(a). The MZMs’ non-Abelian anyonic statistics and their exponential localization at the opposite ends of the nanowire [17] are highly desired properties for designing quantum computation with reduced decoherence issues due to topological protection. The main signature of MZMs in normal metal–superconductor–normal metal junctions is a conductance peak at zero energy [18, 19]. However, the presence of disorder or Andreev bound states in semiconductor–superconductor hybrid devices can cause spurious zero energy modes which are difficult to distinguish from the genuine quasiparticle [20, 21, 22, 23, 24, 25], disrupting the conclusive observation of MZMs thus far. As such, experimental techniques with the sensitivity to make this distinction are sorely needed [26].
Figure 1: (a) The superconducting nanowire in the presence of an external magnetic field and coupled to normal metal leads, $L$ and $R$. The black and red lines denote schematically the exponentially localized probability density of the Majorana zero mode. (b) Schematic procedure of extracting the traversal time from the distance of current cross-correlation peaks.
In this work, we use dynamical current correlations and electron traversal times [27, 28] to identify Majorana zero modes (see Fig. 1(b)). Earlier studies have explored noise in MZM-supporting nanowires [29, 30, 31, 32], and high-frequency noise and cross-correlation measurements enable time-resolved detection in mesoscopic conductors [33, 34, 35, 36, 37, 38, 39, 40]. Motivated by the noninvasive Fano-factor approach [32], we show that the nonequilibrium current cross-correlation yields a parameter-tolerant signature present for topological MZMs and absent for trivial states. To our knowledge, these transient Majorana signatures have not been identified previously.
The nonequilibrium Green’s function methodology [41] is well-suited to the study of electron transport in superconducting nanowires [42, 13, 43, 44, 45, 28, 46]. Recent years have seen the development of various approaches to time-dependent quantum transport based on the Green’s function methodology, including the time-dependent Meir-Wingreen [47, 48, 49] and Landauer-Büttiker (TDLB) approaches [50, 51, 52], as well as the time-linear generalized Kadanoff-Baym ansatz for modeling correlated dynamics [53, 54, 55, 56, 57]. Notably, the TDLB theory is a single-shot approach where time steps can be evaluated independently of each other, which enables trivial parallelization, making it computationally more attractive even compared to the time-linear theory.
The TDLB approach models transient correlations following a partition-free quench process [58], where the leads equilibrate with the nanowire prior to the addition of a bias across the leads at the switch-on time $t_{0}$, and can deal with arbitrary time-dependent biases [59]. It includes the study of time-dependent quantum noise on the signal via calculation of the cross-lead current correlation function [60, 27]
$C_{LR}\left(t+\tau,t\right)\equiv\left\langle\Delta\hat{I}_{L}\left(t+\tau\right)\Delta\hat{I}_{R}\left(t\right)\right\rangle,$ (1)
where $\Delta\hat{I}_{\gamma}\left(t\right)\equiv\hat{I}_{\gamma}\left(t\right)-\left\langle\hat{I}_{\gamma}\left(t\right)\right\rangle$ is the fluctuation of the electronic current in lead $\gamma\in\{L,R\}$. In this expression, $t$ is the observation time and $\tau$ is the relative time. The experimentally-accessible cross correlation between the left and right leads is obtained as the real, lead-symmetrized average
$P(t+\tau,t)\equiv\frac{1}{2}\,\mathrm{Re}\!\left[C_{LR}(t+\tau,t)+C_{RL}(t+\tau,t)\right].$ (2)
Since $\mathrm{Re}[C_{LR}(t+\tau,t)]=\frac{1}{2}\langle\{\Delta\hat{I}_{L}(t+\tau),\Delta\hat{I}_{R}(t)\}\rangle$ [61], $P$ coincides with the Blanter-Büttiker cross correlation in reciprocal, stationary settings ($\mathrm{Re}\,C_{LR}=\mathrm{Re}\,C_{RL}$) [62]. Also in non-reciprocal cases, Eq. (2) provides a lead-exchange symmetric measure of cross correlation. Under stationarity, all quantities depend only on $\tau$, and the associated Fourier transformation, $P(\omega)$, is typically referred to as noise power.
Unlike a dc Fano factor, which time-averages over local and nonlocal channels and can be mimicked by trivial (quasi-)Andreev bound states, transient measurements avoid many of these ambiguities documented in realistic nanowires [63]. In particular, time-resolved cross-lead noise isolates crossed-Andreev conversion and reveals a finite traversal time across the wire, a genuinely nonlocal signature expected when spatially separated Majoranas dominate transport [29, 64]. Consistently, finite-frequency or waiting-time diagnostics predict qualitative differences between topological and trivial cases, reinforcing the advantage of the time domain as a discriminator [65, 66, 67].
In the following, we develop the two-time cross-lead current-current correlator in the TDLB-Nambu framework and show that its time-domain echoes yield a length-linear traversal time containing a nonlocal delay robust to trivial near-zero modes. This builds on Ref. [43], where time-resolved current signatures were studied. The extra contributions to the traversal times are well captured within a novel and simple-to-implement heuristic formula. Finally, we propose an experiment where this intuitive approach can be used to characterize genuine MZMs.
The model.— The non-interacting TDLB picture can be extended to describe the superconducting nanowire in proximity to a bulk $s$-wave superconductor shown in Fig. 1(a). This is a $1D$ nanowire sandwiched between two normal metal conductors, featuring a strong spin-orbit interaction characteristic of, e.g. InSb or Ge. The spin therefore naturally aligns with the $\pm y$-direction, however, a strong external applied $B$-field along the axis of the nanowire breaks time-reversal symmetry and cants the spins along the $\pm z$-direction, introducing a Zeeman splitting $V_{Z}=g\mu_{B}B/2$, where $\mu_{B}$ is the Bohr magneton and $g$ is the Landé factor.
In Bardeen-Cooper-Schrieffer theory, the phonon-mediated attraction between electrons is captured by the pairing field $\Delta$ [68]. This translates to a tractable single-particle picture using a Nambu-spinor representation. We let operators $\hat{d}_{n\sigma}$ and $\hat{d}^{\dagger}_{n\sigma}$ annihilate and create electrons with spin $\sigma$, in spatial location $n$, and introduce a Nambu representation, $\hat{\Phi}_{n}=(\hat{d}_{n\uparrow},\hat{d}_{n\downarrow}^{\dagger},\hat{d}_{n\downarrow},\hat{d}_{n\uparrow}^{\dagger})^{T}$. The Hamiltonian of the isolated wire is given in the Bogoliubov–de Gennes form [69]
$\hat{H}_{w}=\frac{1}{2}\underset{n}{\sum}\left[\hat{\Phi}_{n}^{\dagger}A_{n}\hat{\Phi}_{n}+\left(\hat{\Phi}_{n}^{\dagger}B_{n}\hat{\Phi}_{n}+\text{h.c.}\right)\right],$ (3)
where the on-site contribution is given by
$A_{n}=\left(\begin{smallmatrix}J-\mu+V_{Z}&-\Delta&0&0\\ -\Delta&\mu-J+V_{Z}&0&0\\ 0&0&J-\mu+V_{Z}&\Delta\\ 0&0&\Delta&\mu-J+V_{Z}\end{smallmatrix}\right)$ (4)
and the intra-wire hopping contribution is the matrix
$B_{n}=\left(\begin{smallmatrix}-J/2&0&-\alpha/2&0\\ 0&J/2&0&-\alpha/2\\ \alpha/2&0&-J/2&0\\ 0&\alpha/2&0&J/2\end{smallmatrix}\right),$ (5)
[…]
The energy levels of the leads are subjected to a time-dependent, spatially homogenous shift by a voltage at time $t_{0}$. That is, $\varepsilon_{kL/R}\rightarrow\varepsilon_{kL/R}+V_{L/R}\left(t\right)$ at $t>t_{0}$. The Hamiltonian of the leads is then expanded in the Nambu basis as $\hat{H}_{l}=\frac{1}{2}\underset{k\gamma}{\sum}\hat{\Phi}_{k\gamma}^{\dagger}\tilde{\varepsilon}_{k\gamma}\hat{\Phi}_{k\gamma}$, where $\tilde{\varepsilon}_{k\gamma}\equiv\varepsilon_{k\gamma}\textrm{diag}\left(1,-1,1,-1\right)$ is the energy of state $k$ in the $\gamma$-th lead.
Finally, the lead-nanowire coupling Hamiltonian is given by $\hat{H}_{c}=\frac{1}{2}\underset{nk\gamma}{\sum}\left(\hat{\Phi}_{n}^{\dagger}\tilde{T}_{nk\gamma}\hat{\Phi}_{k\gamma}+h.c\right)$, where $\tilde{T}_{nk\gamma}\equiv T_{nk\gamma}\textrm{diag}\left(1,-1,1,-1\right)$ are the tunneling matrix elements between the $k$-th state of lead $\gamma$ and the $n$-th state of the wire. The Hamiltonian of the total system then reads $\hat{H}_{\mathrm{tot}}=\hat{H}_{w}+\hat{H}_{l}+\hat{H}_{c}$.
The method.— We generalize the TDLB approach to the quantum noise [60] so that all Green’s functions appearing in the current-current correlation function are given in the Nambu representation, where each of the $N$-sites on the nanowire shown in Fig. 1(a) corresponds to four Nambu states. For inter-site spacing $a$ the wire length $L=Na$. The one-particle Green’s function in the Nambu representation is defined as
$\left[\mathbf{G}\left(z_{1},z_{2}\right)\right]_{mn}=-i\left\langle\mathcal{T}_{KP}\left[\hat{\Phi}_{m}\left(z_{1}\right)\otimes\hat{\Phi}_{n}^{\dagger}\left(z_{2}\right)\right]\right\rangle,$ (6)
where $\mathcal{T}_{KP}$ denotes an ordering of times $z_{1}$, $z_{2}$ along the complex Konstantinov-Perel’ contour [70], and the brackets denote a quantum statistical average. We set $\hbar=e=k_{\mathrm{B}}=1$ throughout this work. Quantities denoted in bold are $4N\times 4N$-dimensional matrices. $\mathbf{G}$ satisfies the Kadanoff-Baym integro-differential equations of motion on the contour [71], with integral kernel given by the embedding self-energy $\left[\mathbf{\Sigma}_{em}\left(z_{1},z_{2}\right)\right]_{mn}=\underset{k\gamma}{\sum}\tilde{T}_{mk\gamma}\left[\mathbf{g}_{\gamma\gamma}\left(z_{1},z_{2}\right)\right]_{kk}\tilde{T}_{k\gamma n}$, where $\mathbf{g}_{\gamma\gamma}$ denotes the ‘bare’ Green’s function of lead $\gamma$ and the $\tilde{T}_{mk\gamma}$ are defined in the lead-nanowire coupling $\hat{H}_{c}$. All self-energy components can be expressed in terms of $\mathbf{\Gamma}_{\gamma}=2\pi\sum_{k}\tilde{T}_{mk\gamma}\tilde{T}_{k\gamma n}\delta\left(\varepsilon_{\gamma}^{F}-\tilde{\varepsilon}_{k\gamma}\right)$, where $\varepsilon_{\gamma}^{F}$ is the Fermi energy of lead $\gamma$ [72]. The level-width $\mathbf{\Gamma}=\sum_{\gamma}\mathbf{\Gamma}_{\gamma}$ describes the energy-independent lead-nanowire coupling within the wide-band limit approximation [73]. This approximation is valid because we aim to study low-energy excitations near the Fermi level and the normal metal leads can be safely assumed to have broad energy bands compared to the Majorana energy scale.
The lesser and greater components of the one-particle Green’s function of the nanowire are given by
$\mathbf{G}^{\lessgtr}\left(t_{1},t_{2}\right)=\pm i\int\frac{d\omega}{2\pi}f\left(\pm\omega\right)\underset{\gamma}{\sum}\mathbf{S}_{\gamma}\left(t_{1};\omega\right)\mathbf{\Gamma}_{\gamma}\mathbf{S}_{\gamma}^{\dagger}\left(t_{2};\omega\right),$ (7)
where $f\left(x\right)=\left(e^{\beta\left(x-\mu\right)}+1\right)^{-1}$ is the Fermi-Dirac distribution, $\beta$ is the inverse temperature and $\mu$ is the chemical potential [72].
We have introduced the matrix $\mathbf{S}_{\gamma}\left(t;\omega\right)\equiv e^{-i\mathbf{h}^{\text{eff}}\left(t-t_{0}\right)}\left[\mathbf{G}^{r}\left(\omega\right)-i\int_{t_{0}}^{t}d\bar{t}e^{-i\left(\omega\mathbf{I}-\mathbf{h}^{\text{eff}}\right)\left(\bar{t}-t_{0}\right)}e^{-i\psi_{\gamma}\left(\bar{t},t_{0}\right)}\right]$ in terms of the retarded Green’s function $\mathbf{G}^{r}\left(\omega\right)=\left(\omega\mathbf{I}-\mathbf{h}^{\text{eff}}\right)^{-1}$, where $\mathbf{h}^{\text{eff}}=\mathbf{h}-i\mathbf{\Gamma}/2$ is the effective Hamiltonian of the nanowire considered as an open quantum system, reducing to the ‘bare’ nanowire Hamiltonian $\mathbf{h}$, i.e. expressing $\hat{H}_{w}$ in the Nambu basis, in the decoupling limit $\mathbf{\Gamma}\rightarrow 0$. We also introduced the bias-voltage phase factor $\psi_{\gamma}\left(t_{1},t_{2}\right)\equiv{\int}_{t_{2}}^{t_{1}}d\tau\,V_{\gamma}\left(\tau\right)$. The two-time current correlator may be evaluated as a closed function in the two-time plane:
$C_{LR}\left(t_{1},t_{2}\right)$
$=4\mathrm{Tr}\left\{\mathbf{\Gamma}_{L}\mathbf{G}^{>}\left(t_{1},t_{2}\right)\mathbf{\Gamma}_{R}\mathbf{G}^{<}\left(t_{2},t_{1}\right)\right$.
$\left.+i\mathbf{G}^{>}\left(t_{1},t_{2}\right)\left[\mathbf{\Lambda}_{R}^{+}\left(t_{2},t_{1}\right)\mathbf{\Gamma}_{L}+\mathbf{\Gamma}_{R}\left(\mathbf{\Lambda}_{L}^{+}\right)^{\dagger}\left(t_{1},t_{2}\right)\right]\right$.
$\left.+i\left[\mathbf{\Lambda}_{L}^{-}\left(t_{1},t_{2}\right)\mathbf{\Gamma}_{R}+\mathbf{\Gamma}_{L}\left(\mathbf{\Lambda}_{R}^{-}\right)^{\dagger}\left(t_{2},t_{1}\right)\right]\mathbf{G}^{<}\left(t_{2},t_{1}\right)\right$.
$\left.-\mathbf{\Lambda}_{R}^{+}\left(t_{2},t_{1}\right)\mathbf{\Lambda}_{L}^{-}\left(t_{1},t_{2}\right)-\left(\mathbf{\Lambda}_{L}^{+}\right)^{\dagger}\left(t_{1},t_{2}\right)\left(\mathbf{\Lambda}_{R}^{-}\right)^{\dagger}\left(t_{2},t_{1}\right)\right\}.$ (8)
The trace in Eq. (Current Cross-Correlation Spectroscopy of Majorana Bound States) is taken with respect to the nanowire degrees of freedom. We introduce the lead matrices $\mathbf{\Lambda}^{+}$ and $\mathbf{\Lambda}^{-}$, which correspond to electrons and positively charged holes, respectively, propagating from the leads, as $\mathbf{\Lambda}_{\gamma}^{\pm}\left(t_{1},t_{2}\right)=\pm ie^{-i\psi_{\gamma}\left(t_{1},t_{0}\right)}\int\frac{d\omega}{2\pi}f\left(\pm\omega\right)e^{-i\omega\left(t_{1}-t_{0}\right)}\mathbf{\Gamma}_{\gamma}\mathbf{S}_{\gamma}^{\dagger}\left(t_{2};\omega\right)$ [60]. Physically, the terms in Eq. (Current Cross-Correlation Spectroscopy of Majorana Bound States) are given by lead-lead, lead-nanowire and nanowire-nanowire $e$-$h$ interference effects. Refs. [72] and [60] demonstrate explicitly how the standard Landauer-Büttiker formalisms for the current and current fluctuations, respectively, are recovered from the TDLB formalism in the long-time and static bias limits.
Traversal times.— In the case of the two-terminal system shown in Fig. 1(a), the electronic traversal time $\tau_{\mathrm{tr}}$ can be extracted directly from the lead-symmetrized correlator data, shown in Fig. 1(b), as half of the distance separating the first resonant peaks along the $\tau$-axis. This can be defined more explicitly as the value of the relative delay $\tau$ at which the stationary current cross correlation of Eq. (2) reaches its maximum within a restricted domain $\mathcal{T}$:
$\tau_{\mathrm{tr}}=\underset{\tau\in\mathcal{T}}{\mathrm{arg}\,\mathrm{max}}|P(t+\tau,t)|\quad(t\to\infty).$ (9)
We take the argmax over the restricted domain $\mathcal{T}=[\tau_{\mathrm{min}},\tau_{\mathrm{max}}]$, chosen to exclude trivial near-zero features (before the clean ballistic transit) and later-time reflections [61]. For a simplified model nanowire, $\mu=\alpha=V_{Z}=\Delta=0$, $\tau_{\mathrm{tr}}(k)=L/v_{g}(k)$ with $L=Na$ and the group velocity $v_{g}(k)=d\omega_{k}/dk=Ja\sin(ka)$. The fastest mode occurs at $ka=\pi/2$ leading to $\tau_{\mathrm{min}}=N/J$. The upper bound may be set, e.g., to $\tau_{\mathrm{max}}=2\tau_{\mathrm{min}}$ to isolate the first through-wire peak.
We note that many closely related definitions for traversal times and similar quantum timescales exist in the literature [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. Our definition is intended to ground the traversal time in terms of a quantity that is already accessible to experimentalists, who can construct the cross correlation from repeated non-invasive transport measurements in the leads [87, 88, 89, 90, 37, 91].
Results.— We carry out calculations of the time-dependent cross-correlations for four distinct transport regimes, each specified by a choice of the parameters in our model of a nanowire. The parameters $\mu$, $J$, $V_{Z}$ and $\Delta$ defined above are sufficient to characterize the ordinary and topological superconducting regimes. In addition we study the case of magnetic impurities situated at the wire ends which is characterized by the modified tight-binding parameter $\tilde{V}_{Z}$ described in Ref. [43]. The spurious case of the quasi Majorana mode-supporting nanowire is defined by the usual parameters and an additional smooth confining potential within the nanowire (see Eq. (13) in Ref. [43]). All energies are measured with respect to $J$; times and temperatures are given in terms of $J^{-1}$.
The parameters chosen for our calculations are displayed explicitly in Table 1. In these situations, low-energy states arise from different physical mechanisms with distinct phenomenology. In the trivial superconducting phase, the spectrum is fully gapped and any subgap features correspond to conventional Andreev bound states, which are sensitive to local perturbations [92]. By contrast, MZMs in the topological phase appear as spatially separated, bound, exponentially-decaying end states pinned near zero energy, producing robust zero-bias conductance features and exhibiting nonlocal correlations [93]. Magnetic impurities can induce states occasionally crossing zero energy, but these remain strongly localized and tunable only via local magnetic parameters, lacking topological protection [94]. Also, smooth confinement potentials can produce quasi-Majorana states that mimic many spectroscopic signatures of true MZMs, including apparent zero-bias robustness, yet differ in their strong dependence on confinement geometry and absence of nonlocal parity correlations [95]. Only the low-energy in-gap states are excited by applying a sudden bias voltage $V_{L}(t)=-V_{R}(t)=V_{0}\theta(t)$ with $V_{0}=\Delta/2$.
Parameter sets (in units of $J$)
Model Regime $\mu/J$ $\alpha/J$ $V_{Z}/J$ $\tilde{V}_{Z}/J$ $\Delta/J$
Ordinary phase 0.0 0.5 0.0 N/a 0.1
Topological phase 0.0 0.5 0.25 N/a 0.1
Magnetic impurity 0.0 0.5 0.0 0.67 0.1
Quasi Majorana 2.0 0.5 1.2 N/a 0.1
Table 1: Parameter sets for the different model regimes considered in this work.
We then compute the cross-correlation $P\left(t+\tau,t\right)$ for each parameter set, and present its absolute value normalized by the square of the lead-nanowire coupling strength $\Gamma=0.01$ defined between the two end-sites of the wire and the lead, for each model regime in Fig. 2. Details of the numerical implementation can be found in Appendix B of Ref. [28]. Crucially for our method, the implemented formula is a ‘single-shot’ function in the two-time plane, so there is an identical cost associated with taking the observation time $t$ to be very small or very large. In practice, we set the observation time to $t=5000$, i.e. we work directly in the steady-state regime. We also consider the low-temperature regime (to avoid thermal suppression of topological features) by setting $\beta=200$. With the hopping energies $J$ in sub-eV range for prototypical semiconductor nanowires [96], this corresponds to cryogenic temperatures $\lesssim 10$ K.
Immediately we see by comparing the case of the true topological superconductor (Fig. 2(b)) to the spurious Majorana case (Fig. 2(d)) that there is a qualitative distinction between the corresponding cross-correlation functions. Specifically, in Fig. 2(d) we observe strong, unique resonant peaks at $\tau\simeq\pm 50$, corresponding to the traversal of a coherent wavefront across the nanowire, followed by an oscillatory peak structure caused by internal reflections. This is very similar to the behaviour observed in Ref. [27] for a graphene nanoribbon junction. However, for the genuine topological superconductor in Fig. 2(b), the initial electron traversal event is strongly suppressed with respect to modes effective on a significantly longer timescale.
Figure 2: Stationary current cross-correlation versus relative time shifts $\tau$ for superconducting nanowires of length $N=50$ sites. The nanowire is in the ordinary superconducting phase with no zero modes in panel (a) and it hosts in-gap states at zero energy for panels (b-d). Only the panel (b) hosts the topological Majorana zero mode while panels (c) and (d) host topologically trivial in-gap states. Coupling to the normal-metal leads is specified by the tunneling rate $\Gamma=0.01$, and the temperature is set by $\beta=200$.
Figure 3 shows the traversal times evaluated, in accordance with Eq. (9), for each model regime as a function of the nanowire length $L$. This data clearly supports the argument of a traversal time as the dependence on the wire length is linear. We can roughly decompose the traversal time as $\tau_{\mathrm{tr}}=\tau_{0}+\tau_{\mathrm{cont}}$, where $\tau_{0}$ corresponds to a coherent wavefront traversing the system. The delay time due to the localization close to the contacts $\tau_{\mathrm{cont}}$ is related to the localized edge states and how strongly they hybridize in the presence of leads [61].
These observations can be translated into an heuristic formula for the traversal time if we evaluate it from the sum of the electron dwell times in a site-wise manner, i.e. $\tau_{\mathrm{tr}}=\sum^{N}_{i=1}\tau_{i}$. A site-wise expression for the $\tau_{i}$ (given explicitly in the Supplemental Material [61]) is then evaluated by choosing an effective intra-site coupling $\lambda_{i}$ which follows the inverse of the on-site potential. The resulting heuristic expression is given by
$\tau_{\mathrm{tr}}=\frac{4}{J\pi}\left[\frac{2\left(e^{\delta}-1\right)}{e^{\delta/z}-1}+N-2z\right],$ (10)
where $\delta$ controls the exponential drop-off at each end of the wire, $z$ is the localization length evaluated to the nearest integer, the on-site velocity is numerically estimated to be $v_{i}=\frac{\pi}{2}\lambda_{i}$ and $N=L/a$ is the wire length. Note that this expression scales linearly with $N$, as expected. The estimated traversal times $\tau_{\mathrm{tr}}\sim 100$ in units of inverse hopping would correspond to a temporal resolution of picoseconds, for prototypical semiconductor nanowires.
Figure 3: Extracted traversal times in terms of the nanowire length $L$. Simple linear fits (dashed lines) are performed as $\tau_{\mathrm{tr}}=\tau_{\mathrm{cont}}+L/v$, where $\tau_{\mathrm{cont}}$ and $1/v$ are obtained as the intercept and slope, respectively. For the Majorana zero mode, a fit with Eq. (10) is also included (solid black line) with the parameters $\delta=1.402$ and $z=4$. Other model parameters are $\Gamma=0.01$ and $\beta=200$.
Conclusions.— In this work we analyzed time-resolved cross-lead current correlations in proximitized nanowire junctions within the TDLB framework, using a lead-symmetrized correlator to operationally define and discern an electronic traversal time from transient noise. We evaluated this diagnostic across conventional superconducting, topological Majorana, magnetic-impurity, and quasi-Majorana regimes, finding a robust linear growth of traversal time with wire length and regime-specific temporal fingerprints that distinguish genuine MZMs from spurious near-zero modes. We further distilled a compact heuristic capturing both length scaling and end-localization effects, and verified its quantitative accuracy against simulations. Finally, the predicted picosecond temporal scale expected for prototypical semiconductor nanowires indicates that traversal-time echoes should be observable with current experimental platforms.
To further estimate the time scales relevant for experimental observation, Mourik et al. [4] reported an induced superconducting energy gap of about $250$ µeV in a hybrid superconductor-semiconductor nanowire device. This energy scale corresponds to characteristic in-gap oscillations with a period of roughly $16$ ps, or equivalently a frequency of about $60$ GHz. On the other hand, McIver et al. [97] have demonstrated high-frequency transport measurements using photoconductive switches capable of sub-picosecond temporal resolution. Our framework, employing a noninvasive current cross-correlation probe to identify and distinguish such in-gap states, would therefore call for a stroboscopic implementation of a two-lead setup, where time-resolved currents are detected at both contacts. In this configuration, the nanowire could also be driven out of equilibrium optically rather than by the voltage-bias quench considered in this work.
Future studies could explore Floquet Majorana bound states [98]. These can be implemented within the TDLB framework via the incorporation of a periodic driving voltage, offering an alternative way to manipulate and control topological states. Another interesting possibility is to apply the same time-resolved cross-lead-correlator diagnostic to platforms hosting $\mathbb{Z}_{n}$ parafermionic zero modes [99]. In such systems, fractionalized quasiparticles with effective charge $e^{\ast}{=}2e/n$ and the associated $4\pi/n$-Josephson periodicity should imprint distinct traversal-time fingerprints, e.g., multiple delayed branches and fractional noise scaling, providing a clear discriminant from Majorana (n=2n{=}2) responses and a route to time-domain identification of non-Abelian orders beyond MZMs.
Acknowledgments.— M. R. and E. C. acknowledge support from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101178170 and by the Israel Science Foundation under grant agreement No. 2208/24. C. F. acknowledges the financial support from the Research Council of Finland through the Finnish Centre of Excellence in Quantum Technology (Project No. 352925) and the Finnish Quantum Flagship (Project No. 358877). R. T. acknowledges the financial support of the Jane and Aatos Erkko Foundation (Project EffQSim) and the Research Council of Finland through the Finnish Quantum Flagship (Project No. 359240). We also acknowledge CSC—IT Center for Science, Finland, for computational resources.